# Going Forward-Forward in Distributed Deep Learning

### v1

@@ Abstract
This paper introduces a new approach in distributed deep learning, utilizing Geoffrey Hinton’s Forward-Forward (FF) algorithm to enhance the training of neural networks in distributed computing environments. Unlike traditional methods that rely on forward and backward passes, the FF algorithm employs a dual forward pass strategy, significantly diverging from the conventional backpropagation process. This novel method aligns more closely with the human brain’s processing mechanisms, potentially offering a more efficient and biologically plausible approach to neural network training. Our research explores the implementation of the FF algorithm in distributed settings, focusing on its capability to facilitate parallel training of neural network layers. This parallelism aims to reduce training times and resource consumption, thereby addressing some of the inherent challenges in current distributed deep learning systems. By analyzing the effectiveness of the FF algorithm in distributed computing, we aim to demonstrate its potential as a transformative tool in distributed deep learning systems, offering improvements in training efficiency. The integration of the FF algorithm into distributed deep learning represents a significant step forward in the field, potentially revolutionizing the way neural networks are trained in distributed environments.
@@ 1 Introduction
Training deep neural networks, often consisting of over hundreds of layers, is a very time consuming process that can take several weeks, using the well-known backpropagation [1] as the standard learning algorithm. While this process works well in a single computational unit, parallelization of these models presents significant challenges due to the sequential nature of the backpropagation learning algorithm, as illustrated in Figure 1. In backpropagation, the gradient of each layer depends on the gradient calculated at the subsequent layer. In a distributed setting, this sequential dependency necessitates extensive waiting times between nodes, as each node must wait for gradient information from its successor before it can proceed with its calculations. Moreover, the need for constant communication between nodes to transfer gradient and weight information can lead to significant communication overhead. This is particularly problematic in large-scale neural networks where the volume of data to be transferred can be substantial.
@@
Figure 1: Challenges in parallelizing the traditional backpropagation model. F denotes the forward pass, first parameter of F is the layer that the pass is made, and the second parameter is the data that is passed through that layer. B denotes the backward pass and takes the prediction and the real label for calculating the gradient.
@@
Figure 2: Parallelization using the FF algorithm. G is the Goodness (to be defined) of the Forward pass.
This paper explores the integration of the FF algorithm into distributed deep learning systems, particularly focusing on its potential to enhance the efficiency and effectiveness of training large-scale neural networks. We demonstrate the effectiveness of the Pipeline Forward-Forward (PFF) algorithm, which reduces the training duration by almost 5-fold for a 5-layer network while preserving the accuracy of the original algorithm.
@@ 2 Literature Review
Distributed and parallel neural networks have attracted widespread interest and investigation from researchers, driven by the promise of enhanced computational efficiency and scalability in handling large-scale data and complex models. We used several comprehensive surveys from 2019 [6] and 2022 [7]. These surveys explain the parallelism methods including data, model and pipeline that we employ in our methods.
PipeDream [3] is a deep neural network training system using backpropagation that efficiently combines data, pipeline, and model parallelism by dividing the layers of the model into multiple stages that contains multiple consecutive layers. Unlike our method, PipeDream inherits the limitations of the classic back propagation algorithm.
Similar to PipeDream, GPipe [2] is a distributed deep learning library that updates the weights synchronously during backward pass, which differentiates it from PipeDream. However, it also suffers from additional dependencies introduced by the classic back propagation algorithm.
@@ 3 Forward-Forward Algorithm
The FF algorithm, proposed by Geoffrey Hinton [5], presents a novel approach to training neural networks. Inspired by the biological processes of the brain, the FF algorithm eliminates backpropagation and utilizes two forward passes to update network weights. Those passes are called respectively, the positive pass and the negative pass.
The positive pass strengthens the recognition of actual data patterns, while the negative pass weakens the recognition for patterns do not match the target data. While other approaches are possible, we will follow Hinton‘s approach where a positive sample is generated by adding the correct label on the top-left corner of a sample image from the training dataset, whereas a negative sample is one where a wrong label is selected (one that does not match the image class).
The most important feature of the FF algorithm for this paper is that its training process doesn’t involve a global backward pass, eliminating the need for backpropagation, which is not deemed as biologically plausible and carries challenges for training deep networks.
Training.
The aim of the learning is to make the ”goodness value” of real (positive) data higher than some threshold and lower it for negative data. The goodness of a layer is simply the sum of the squares of the activities of the rectified linear neurons in that layer. This is then related to the probability that an input vector is positive (i.e., real):
$p(real)=\sigma(\Sigma_{j}y_{j}^{2}-\theta)$
@@
During inference, ’goodness’ score is aggregated across all but the first hidden layer of the network for each label. Then, the label resulting in the maximum goodness is selected as the predicted class.
An alternative inference method was suggested where the input image is labelled with a neutral label (since we do not know the correct class) and a Softmax layer is trained using the activity vectors collected from all layers except the first hidden layer [5].
Figure 3: Traditional Forward-Forward algorithm with three layers and three splits (S = 3).
@@ 4 PipelineFF: A Pipeline Forward-Forward Algorithm
In this section we propose three Pipeline Forward Forward (PFF) algorithms: Single Layer PFF (4.1), All Layers PFF (4.2), and Federated PFF (4.3).
The Single Layer PFF employs an architecture where each node is dedicated to training only one layer, while All Layers and Federated PFF algorithms distribute the training load more evenly, by having every node train all network layers in turn. All models enjoy model and pipeline parallelism, while the federated PPF also takes advantage of data parallelism.
In all of the proposed distributed FF models, we split the total number of epochs $E$ into the total number of the splits or chapters $S$ to facilitate distributed training. Each chapter thus takes $C=E/S$ training epochs. Consequently, each client will process $S$ chapters to complete the training for $E$ epochs. The variable $N$ denotes the number of nodes in the distributed system. The input dataset is abstracted as $x$, which is a concatenated form of the positive and negative datasets, simplifying the explanation. $L_{i}(x)$ is used to denote the output of the $i$-th layer of the neural network after a forward pass with the input data $x$. Finally, $Train(L_{i},x,n)$ is an abstracted function that trains layer $L_{i}$ using the data $x$ for $n$ epochs.
4.1 Single Layer PFF Algorithm
Figure 4: Single Layer Pipeline Forward-Forward Algorithm Example with three layers and three splits.
Single Layer PFF employs an architecture where each node is dedicated to training only one layer, as shown in Fig. 4. It reduces data transfer between nodes.
The distributed training example provided in Figure 4 illustrates a scenario with three compute nodes and a three-layer neural network. The number of nodes, $N$, is selected to be equal to the number of layers $L$ and each node trains the corresponding layer of the network. Already with this simple architecture, we achieve parallelism among the nodes, unlike the purely sequential FF algorithm. This procedure ensures that while one node trains a layer, the previous node can simultaneously train the preceding layer. Below, we detail the algorithm with a pseudo code and the training process step-by-step.
@@ 4.1 Single Layer PFF Algorithm
Iteration and Convergence: The process iterates for $S$ chapters, with each node refining its layer based on the progressively improved outputs from the preceding layer. Convergence is achieved when all layers have been trained for $E$ epochs across all splits.
Algorithm 1 Single Layer Training Procedure
1:for each $chapter$ in $range(split)$ do
@@
3:     for each $layerIndex$ in $range(myIndex)$ do
@@
6:     for each $miniEpoch$ in $range(epochs/split)$ do
@@
9:              $trainLayer(myIndex,batch)$
@@
4.2 All Layers PFF Algorithm
Figure 5: All Layers Pipeline Forward-Forward Algorithm Example with three layers and six splits.
The All Layers PFF model builds upon the Single Layer approach by allowing each compute node to train the entire network, affording a more balanced load among the nodes.
@@ 4.2 All Layers PFF Algorithm
Algorithm 2 All Layers Training Procedure
1:for each $chapter$ in $range(split)$ do
@@
3:     for each $layerIndex$ in $range(layerAmount)$ do
@@
6:         for each $miniEpoch$ in $range(epochs/split)$ do
@@ 4.3 Federated PFF Algorithm
In contrast to the All Layers PFF Algorithm, the Federated PFF Algorithm employs the same training procedure but differentiates in the data it utilizes for training. Each node trains on its own local dataset rather than a shared one. This preserves data privacy and autonomy as there is no need to share raw data with other nodes or a centralized server. While the data remains localized, nodes exchange model updates.
This allows nodes to benefit from learning carried out by their counterparts, which iteratively improves the global model that benefits from diverse data without central aggregation.
This federated structure is not only crucial for scenarios requiring data privacy but also enables distributed training over networks where central data aggregation is impractical or undesired.
@@ 5 Experimental Evaluation
We obtain and compare several models by modifying the distributed model (Single Layer or All Layers), negative data selection process (Adaptive, Fixed or Random) and classification strategies (Goodness and Softmax).
For selecting negative samples, we implemented Adaptive-NEG, the method proposed in [5], which selects the most predicted incorrect label as the negative label. The second approach is Fixed-NEG, which involves choosing random incorrect labels for each instance in the beginning of the training. The third approach, called Random-NEG, selects the random incorrect labels at the end of each chapter. The last two are evaluated to speed up the training.
As for classifying approach, we adopt the two approaches proposed in [5]. In the Goodness based approach, a sample is classified into the class that results in the highest goodness score (considering all but the first layer). In the Softmax approach, a single layer classification head is added to the network with the softmax activation.
Each of the models are implemented with two PFF variants: Single Layer model (Section  4.1) and All Layers model (Section  4.2). The Sequential PFF, on the other hand, is equivalent to the original FF algorithm, where the node amount $N=1$ for comparing the models using the same code base.
Our chosen architecture for the experiment consists of a neural network with a configuration of [784, 2000, 2000, 2000, 2000] where the numbers indicate the number of nodes in each layer. Thus the input, which is 734-dimensional, is followed by four hidden layers with 2000 nodes each. Each of these layers employs the ReLU activation function and the threshold coefficient is set to 0.01, as in [5]. The network is trained and tested using the MNIST dataset, using 60,000 training instances with mini-batches of size 64 and testing on a separate test set comprising 10,000 instances for 100 epochs and 100 splits. The Adam optimizer is used for both the FF Layers and the Softmax layer (trained using Backpropagation) is set initially at 0.01 and 0.0001 respectively and cooldowns after the 50th epoch.
@@ 5.1 Comparison of FF, DFF and PFF Models
While the results of the compared models are obtained from their respective publications, all our models are trained for the same total number of epochs. The Adaptive-NEG Goodness model with Sequential implementation is basically the same as [5] and it indeed achieves almost the same accuracy as Hinton‘s Matlab implementation [11]. It enables us to compare the training time of the model with our Python implementation.
Among negative data selection methods, the Adaptive-NEG approach performs best, as expected, since it selects the negative data at each chapter. Among the distributive alternatives, the All Layers implementation achieves near top results with over three-fold improvement in training time (2,980 versus 11,190 seconds). Moreover, in the Single Layer method, negative samples in Adaptive-NEG are only generated and published to other nodes from the last node. In contrast, in the All Layers method, each node independently computes the negative samples at the end of each chapter. This way, the tasks are distributed better, concluding to a faster training. Furthermore, to increase time efficiency, in the Single Layer implementation other nodes do not wait for the generated negative samples from the last node at the end of each chapter. Indeed, other nodes begin using the updated negative data immediately upon receiving it. This characteristic slightly reduces the accuracy of the Single Layer model.
@@
Model Implementation Time (s) Accuracy (%)
@@
Adaptive-NEG Goodness Single Layer 5,254.87 98.43
All Layers 2,980.76 98.51
@@
Random-NEG Goodness Single Layer 1,974.10 98.26
All Layers 2,008.25 98.17
@@
Fixed-NEG Goodness Single Layer 1,920.80 97.94
All Layers 1,978.21 97.89
Notably, the Random-NEG model achieves a level of accuracy comparable to the Adaptive-NEG model (e.g. 98.17 vs 98.51 accuracy for the All Layers model), while having shorter training times (2008 vs 2980 seconds). Finally, Adaptive-NEG and Random-Neg models performs better than Fixed-NEG model in every implementation, as the variability in the selection of the negative data enhances the robustness of the system.
5.2 Classifier Model Comparison
In the first evaluation, we compared PFF variants that used the Goodness approach for classification and established the Adaptive-NEG Goodness model as the proposed baseline. The second experiment compares the performance of this model with the one using the Softmax approach. As mentioned before, the goodness method is the first method provided in [5] and uses only the FF layers. Softmax method is presented as an alternative, predicting the output using a Softmax layer that gets the activations of the hidden FF layers except the first.
The corresponding results can be found in Table 2. The results show that the Adaptive-NEG Softmax model (results from Table 1) trains faster than the Adaptive-NEG Goodness model, for all different implementations. This accelerated training is due to the single-step prediction in the softmax approach, rather than predicting across 10 different classes. However, since the predictions made by the Softmax layer do not directly indicate the FF Network’s overall performance, we observe a slight drop in accuracy.
Table 2: Softmax Analysis of Adaptive-NEG. The first three row results are from Table 1.
Model Implementation Time (s) Accuracy (%)

### v2

@@ Abstract
We introduce a new approach in distributed deep learning, utilizing Geoffrey Hinton’s Forward-Forward (FF) algorithm to speed up the training of neural networks in distributed computing environments. Unlike traditional methods that rely on forward and backward passes, the FF algorithm employs a dual forward pass strategy, significantly diverging from the conventional backpropagation process. This novel method aligns more closely with the human brain’s processing mechanisms, potentially offering a more efficient and biologically plausible approach to neural network training. Our research explores different implementations of the FF algorithm in distributed settings, to explore its capacity for parallelization. While the original FF algorithm focused on its ability to match the performance of the backpropagation algorithm, the parallelism aims to reduce training times and resource consumption, thereby addressing the long training times associated with the training of deep neural networks. Our evaluation shows a 3.75 times speed up on MNIST dataset without compromising accuracy when training a four-layer network with four compute nodes. The integration of the FF algorithm into distributed deep learning represents a significant step forward in the field, potentially revolutionizing the way neural networks are trained in distributed environments.
@@ 1 Introduction
Training deep neural networks, often consisting of over hundreds of layers, is a very time consuming process that can take several weeks, using the well-known backpropagation [1] as the learning algorithm. On the other hand, parallelization of this algorithm presents significant challenges due to the sequential nature of the backpropagation learning algorithm, as illustrated in Figure 1.
In backpropagation, the gradient of each layer depends on the gradient calculated at the next layer. In a distributed setting, this sequential dependency necessitates extensive waiting times between computing nodes, as each node must wait for the gradient information to backpropagate from its successor before it can proceed with its calculations. Moreover, the need for constant communication between nodes to transfer gradient and weight information can lead to significant communication overhead. This is particularly problematic in large-scale neural networks where the volume of data to be transferred can be substantial.
@@
Figure 1: Challenges in parallelizing traditional backpropagation involve managing the forward (F) and backward (B) passes, both applied to the layer specified by their first parameter.
@@
Figure 2: Parallelization using the FF algorithm.
This paper explores the integration of the FF algorithm into distributed deep learning systems, particularly focusing on its potential to improve the efficiency and effectiveness of training large-scale neural networks. We demonstrate the effectiveness of the Pipeline Forward-Forward (PFF) algorithm, which reduces the training duration by almost 5-fold for a 5-layer network while preserving the accuracy of the original algorithm.
@@ 2 Literature Review
Distributed and parallel neural networks have attracted widespread interest and investigation from researchers, driven by the promise of improved computational efficiency and scalability in handling large-scale data and complex models. We used several comprehensive surveys from 2019 [6] and 2022 [7]. These surveys explain the parallelism methods including data, model and pipeline that we employ in our methods.
PipeDream [3] is a deep neural network training system using backpropagation that efficiently combines data, pipeline, and model parallelism by dividing the layers of the model into multiple stages that contains multiple consecutive layers. Unlike our method, PipeDream inherits the limitations of the classic back propagation algorithm. Similar to PipeDream, GPipe [2] is a distributed deep learning library that updates the weights synchronously during backward pass, which differentiates it from PipeDream. However, it also suffers from additional dependencies introduced by the classic back propagation algorithm.
Besides the above methods, there are several proposed methods for efficient deep learning training that attempts to overcome the limitations of back propagation by proposing alternatives to back propagation. One of them is called local parallelism [8]. This method divides the layers into number of blocks, and use this block’s local loss to adjust the parameters of the layers that are within that block. Even though this method achieves a significant speed up compared to back propagation, it is effective to an extent because the local loss of the blocks are not proportional to global loss.
@@ 3 Forward-Forward Algorithm
The FF algorithm, proposed by Geoffrey Hinton [5], presents a novel approach to training neural networks that doesn’t involve backpropagation, which is the standard learning algorithm used in training neural networks.
Inspired by the biological processes of the brain, the FF algorithm trains layers individually and sequentially, utilizing two forward passes to update the layer weights. Those passes are called respectively, the positive pass and the negative pass. The positive pass adjusts the weights to increase the ”goodness” of the positive or real data, while the negative pass does the opposite for negative data, as defined below.
Negative Data.
In [5], positive and negative samples are generated based on the existence of a black border in the MNIST dataset which is used in evaluations. A positive sample is generated by adding the correct label on the top-left corner of the input image (a 10-pixel area in the border denotes the label where a 1-of-C encoding is used), whereas a negative sample is one where a wrong label is selected. Other alternatives for creating negative data, including unsupervised methods, are also possible and explored in [5].
Goodness Function.
The goodness of a layer explored in [5] is the sum of the squares of the activities of the rectified linear neurons in that layer and the aim of the learning is to make it above some threshold value for positive (real) data and below the threshold for negative data. This can be captured by applying the logistic function to the sum of squared activities $y_{j}^{2}$ minus a threshold $\theta$:
$p(real)=\sigma(\Sigma_{j}y_{j}^{2}-\theta)$ (1)
@@
Two methods have been proposed for predicting the label of an input image. In the Goodness prediction, the input is run with 10 different labels where the target vector of a single 1 is placed in the correct class is indicated and the activations from all but the first hidden layer of the network are accumulated. Then, the label resulting in the maximum goodness is selected as the predicted class.
In the Softmax prediction, the unlabelled input image is labelled with a neutral label that consists of 0.1 values for each label and the activity vectors collected from all but the first hidden layer are fed into a Softmax layer which is trained to predict the class. The softmax layer is trained using backpropagation either at the end of the training or during the training. To remark since the softmax layer does not propagate to the FF layers, it can also be trained separately like a FF layer. This prediction approach is faster for inference as there is only a single pass; on the other hand, it is sub-optimal to the Goodness prediction which better corresponds to how the network was trained.
Sequential Training of the Forward-Forward Algorithm.
The layers of the network are trained for several epochs where an epoch means passing the whole training data from the network once. Since FF does not involve backpropagation, layers can be trained independently, each layer can be trained for more than one epoch waiting the upper layers to finish the previous epochs. For instance, assuming that the network will be trained for 60 epochs, we can train the first layer for 30 epochs, and use this half-trained layer-1 for training the second layer for its 30 epochs and so on.
Moreover, training a layer for total number of epochs and providing its output to the next layer is not good for accuracy. Training each layer for small number of epochs allow fine grained adjustments on this layer and yields better results. So, we divide training process of each layer into splits, basically a continuous number of undivided training number of epochs for this layer.
Figure 3: Comparison of the Forward-Forward algorithm with split 1 and split 100 (S = 1, 100) across three layers.
@@ 4 PipelineFF: A Pipeline Forward-Forward Algorithm
In this section we propose three Pipeline Forward Forward (PFF) algorithms: Single-Layer PFF (4.1), All-Layers PFF (4.2), Federated PFF (4.3) and Performance-Optimized PFF (4.4).
The Single-Layer PFF employs an architecture where each node is dedicated to training only one layer, while All-Layers and Federated PFF algorithms distribute the training load more evenly, by having every node train all network layers in turn. All models enjoy model and pipeline parallelism, while the federated PFF also takes advantage of data parallelism.
In all of the proposed pipelined FF models, we split the total number of epochs $E$ into the total number of the splits or chapters $S$ to facilitate distributed training. Each chapter thus takes $C=E/S$ training epochs. Consequently, each client will process $S$ chapters to complete the training for $E$ epochs. The variable $N$ denotes the number of nodes in the distributed system. The input dataset is abstracted as $x$, which is a concatenated form of the positive and negative datasets, simplifying the explanation. $L_{i}(x)$ is used to denote the output of the $i$-th layer of the neural network after a forward pass with the input data $x$. Finally, $Train(L_{i},n)$ is an abstracted function that trains layer $L_{i}$ for $n$ epochs.
4.1 Single-Layer PFF Algorithm
Figure 4: Single-layer Pipeline Forward-Forward Algorithm example with three layers and three splits.
Single-Layer PFF employs an architecture where each node is dedicated to training only one layer, as shown in Fig. 4. This procedure ensures that while one node trains a layer, the previous node can simultaneously train the preceding layer. Already with this simple architecture, we achieve parallelism among the nodes, unlike the sequential FF algorithm.
The distributed training example provided in Figure 4 illustrates a scenario with three compute nodes and a three-layer neural network. The number of nodes, $N$, is selected to be equal to the number of layers $L$ and each node trains the corresponding layer of the network. Below, we detail the algorithm with a pseudo code and the training process.
@@ 4.1 Single-Layer PFF Algorithm
Iteration and Convergence: The process iterates for $S$ chapters, with each node refining its layer based on the progressively improved outputs from the preceding layer. Convergence is achieved when All-Layers have been trained for $E$ epochs across all splits.
Algorithm 1 Single-Layer Training Procedure
1:for $chapter$ = $1$ to $split$ do
@@
3:     for $layerIndex$ = $1$ to $currentClientIndex$ do
@@
6:     for $miniEpoch$ = $1$ to $epochs/split$ do
@@
9:              $trainLayer(currentClientIndex,batch)$
@@
4.2 All-Layers PFF Algorithm
Figure 5: All-Layers Pipeline Forward-Forward Algorithm example with three layers and six splits.
The All-Layers PFF model builds upon the Single-Layer approach by allowing each compute node to train the entire network, affording a more balanced load among the nodes.
@@ 4.2 All-Layers PFF Algorithm
Algorithm 2 All-Layers Training Procedure
1:for $chapter$ = $1$ to $split$ do
@@
3:     for $layerIndex$ = $1$ to $numLayers$ do
@@
6:         for $miniEpoch$ = $1$ to $epochs/split$ do
@@ 4.3 Federated PFF Algorithm
In contrast to the All-Layers PFF Algorithm, the Federated PFF Algorithm employs the same training procedure but differentiates in the data it utilizes for training. Each node trains on its own local dataset rather than a shared one. This preserves data privacy and autonomy as there is no need to share raw data with other nodes or a centralized server. While the data remains localized, nodes exchange model updates. This allows nodes to benefit from learning carried out by their counterparts, which iteratively improves the global model that benefits from diverse data without central aggregation. This federated structure is not only crucial for scenarios requiring data privacy but also enables distributed training over networks where central data aggregation is impractical or undesired.
Figure 6: Federated Forward-Forward Algorithm example with 3 layers and 6 split.
4.4 PFF with A New Goodness Function
The Forward-Forward algorithm [5] states that different goodness functions may be used while training each layer individually. In this section, we propose a new PFF algorithm that uses a goodness function based on classification accuracy. This model is called Performance-Optimized PFF.
Specifically, we add a softmax layer to the network layer that is trained with the PFF algorithm and update that layer’s weights using backpropagation for only the two layers (the added softmax and the actual layer), as illustrated in Figure 7. In other words, we train each layer to maximize the classification accuracy with the help of an added softmax layer; hence, this approach is only suitable for supervised learning problems. Furthermore there is no negative data, but the rest of the training is done exactly in accordance with the FF algorithm.
Figure 7: Forward and Backward passes of original FF
Figure 8: Forward and Backward passes of the proposed FF algorithm, with a separate classifier block for each layer.
@@ 5 Experimental Evaluation
We obtain and compare several models by modifying the distributed model (Single-Layer or All-Layers), negative data selection (Adaptive, Fixed or Random), classification strategies (Goodness and Softmax) and goodness function.
There are two PFF variants: Single-Layer and All-Layers models, as explained in Section  4.1 and Section  4.2 respectively. These variants are compared to the sequential implementation (node count $N=1$), which is equivalent to the original FF algorithm and can be used for comparing the models using the same code base.
For selecting negative samples, we implemented AdaptiveNEG, the method proposed in [5], which selects the most predicted incorrect label as the negative label. The second approach is FixedNEG, which involves choosing random incorrect labels for each instance in the beginning of the training. The third approach, called RandomNEG, selects the random incorrect labels at the end of each chapter. The last two are proposed to speed up the training.
For the classification approach, we implemented both approaches proposed in [5]. In the Goodness based approach, a sample is classified into the class that results in the highest goodness score (considering all but the first layer). In the Softmax approach, a single layer classification head is added to the network with the softmax activation.
Finally, we evaluated a new goodness function for supervised learning scenarios.
5.1 Implementation Details
Our chosen architecture is the same feed-forward neural network as in [5]. Its configuration is [784, 2000, 2000, 2000, 2000], where the numbers indicate the number of nodes in each layer. Thus the input, which is 784-dimensional, is followed by four hidden layers with 2000 nodes each. Each of these layers employs the ReLU activation function.
The network is trained and tested using the MNIST dataset, using 60,000 training instances with mini-batches of size 64 and testing on a separate test set comprising 10,000 instances for 100 epochs and 100 splits. The Adam optimizer is used for both the FF Layers and the Softmax layer (trained using Backpropagation) is set initially at 0.01 and 0.0001 respectively and cooldowns after the 50th epoch. The threshold coefficient $\theta$ in Eq. 1 is set to 0.01, as in [5].
@@ 5.2 Comparison of FF, DFF and PFF Models
The AdaptiveNEG approach, which selects negative data based on the network’s performance at each chapter, outperforms both RandomNEG and FixedNEG models in terms of accuracy. Remarkably, despite shorter training times (7,178 vs 11,190 seconds), RandomNEG performs close to AdaptiveNEG in terms of accuracy (98.33 vs 98.52 in Sequential), making these models more accurate than FixedNEG in every implementation.
The AdaptiveNEG Goodness model with Sequential implementation is basically the same as [5] and it indeed achieves almost the same accuracy as Hinton‘s Matlab implementation [12]. It enables us to compare the training time of the model with our Python implementation.
Furthermore, AdaptiveNEG with All-Layers implementation matches the top accuracy of its sequential version (98.51% vs. 98.52%) but in about a quarter of the time (2,980 vs. 11,190 seconds), using four compute nodes. This finding showcases the PFF Algorithm’s efficiency in distributing the training, achieving 94% utilization (3.75/4), significantly speeding up the process. The All-Layers method speeds up training for AdaptiveNEG by efficiently distributing tasks between nodes, allowing each to compute its own negative samples after every chapter, unlike the Single-Layer approach where the last node generates and publishes the generated labels.
@@
Model Implementation Training Time (s) Test Accuracy (%)
@@
AdaptiveNEG-Goodness Single-Layer 5,254.87 98.43
All-Layers 2,980.76 98.51
@@
RandomNEG-Goodness Single-Layer 1,974.10 98.26
All-Layers 2,008.25 98.17
@@
FixedNEG-Goodness Single-Layer 1,920.80 97.94
All-Layers 1,978.21 97.89
5.3 Classifier Mode Comparison for AdaptiveNEG
In the first evaluation, we compared PFF variants that used the Goodness approach for classification and established the AdaptiveNEG Goodness model as the proposed baseline. The second experiment compares the performance of this model with the one using the Softmax approach. As mentioned before, the goodness method is the first method provided in [5] and uses only the FF layers. The Softmax method on the other hand is presented as a faster alternative, predicting the output using a Softmax layer that gets the activations of the hidden FF layers (except the first). The corresponding results can be found in Table 2.
The results show that the AdaptiveNEG Softmax model trains faster than the AdaptiveNEG-Goodness model, for all different implementations, with time savings ranging from 25-53% (e.g. 1886.42 versus 2980.76 secs). This accelerated training is due to the single-step prediction in the softmax approach, rather than predicting across 10 different classes. However, the accuracies all slightly decrease compared to comparable Goodness models.
Table 2: Classification mode comparison for AdaptiveNEG. The first three row results indicated in italic denote the baseline model from Table 1. Bold results indicate suggested models for each classification mode.
Model Implementation Training Time (s) Test Accuracy (%)